OpenStudy (anonymous):

Write in simplest form 1/4 + 5/6 + 3/8

OpenStudy (anonymous):

first you find the common denominator... which would be 24... so now you make it so that every fraction has that denominator... giving you..... \[\frac{6}{24} + \frac{20}{24} + \frac{9}{24}\]

OpenStudy (anonymous):

now just add those all up and divide by the GCF and there's your simplest form!

OpenStudy (anonymous):

which means you add 6 and 20 and 9.... so you get..... 35

OpenStudy (anonymous):

lol thank you

OpenStudy (anonymous):

anytime.... but do note that its 35/24

OpenStudy (anonymous):

Find the difference: 11/w - 5/w (2 points) do you know this one?